How many apples are on each tray if man carrying a square tray dropped his onto floor and girl carrying 4 trays drops 1 row of apples off each tray and there are 77 on the floor?
Is it ax^2+bx+c=0? If so, is 1(x)2+4(x)-77=0? How do I get the number representing how many apples are on each tray? Thank you,

Answer by ptaylor(2198)

.Is it ax^2+bx+c=0? If so, is 1(x)2+4(x)-77=0? How do I get the number representing how many apples are on each tray? Thank you,
OK!!! GOOD WORK!!!! ALL YOU NEEDED TO DO WAS SOLVE THE QUADRATIC BUT IT MAY WORK BETTER IF YOU SET THE PROBLEM UP SIMILAR TO THE MANNER THAT I DID BELOW. IN OTHER WORDS, SPECIFY WHAT THE UNKNOWNS REPRESENT
Let x=number of apples that are in each row on each tray
Then total number of apples on each tray=x^2 (Since square trays are specified, the number of apples in the rows are the same as the number of apples in the columns)
Now the man drops x^2 apples
and the girl drops 4x apples
So the total number of apples on the floor is
x^2+4x and we are told that this equals 77
Thus, our equation to solve is:
x^2+4x=77 or
x^2+4x-77=0------quadratic in standard form and it can be factored:
(x+11)(x-7)=0
x=-11---------------no good!!! can't have negative apples
x=7----------------number of apples that are in each row on each tray
x^2=49-------------number of apples on each tray
CK
man drops 49
girl drops 7*4=28
49+28=77
77=77

Hope this helps---ptaylor